If the intensity of one sound is 7.24 x 10-4 W/m2 and the intensity of another sound is 6.31 x 10-3   W/m2, what is the difference of their sound intensity levels?


1
Expert's answer
2021-02-17T11:00:15-0500

By definition, the intensity levels of both sounds are (respectively):


"L_1 = 10\lg\left( \dfrac{I_1}{I_0} \right) \space dB\\
L_2 = 10\lg\left( \dfrac{I_2}{I_0} \right)\space dB"

where "I_1 = 7.24\times 10^{-4}W/m^2" is the inensity of the first sound, "I_2 = 6.31\times 10^{-3}W/m^2" is the intensity of the second sound, and "I_0" is some basic intensity.

The difference in levels is:


"L_2-L_1 = 10\lg\left( \dfrac{I_2}{I_0}\right) - 10\lg\left( \dfrac{I_1}{I_0}\right) = 10\lg\left( \dfrac{I_2}{I_1}\right) \\
L_2-L_1 = 10\lg\left( \dfrac{6.31\times 10^{-3}}{7.24\times 10^{-4}}\right) \approx 9.4\space dB"

Answer. "9.4\space dB"